Austin Enterprises makes and sells two types of phones. Management is trying to determine the most profitable mix. Sales prices, demand and use of manufacturing inputs follows:

Sales Price:

Standard Premium
$400 $540

Input requirements per unit:

Direct Material $20 $40
Direct labor (hours) 1.2 2.0
Machine Hours 1.5 1.0

Costs:

Variable Costs

Direct Labor (cost per hour) 25
Factory Overhead(per MH) 3
Marketing (percent of sales price) 10%

Fixed Costs:

Manufacturing $230,000
Marketing 80,000
Administration 120,000

The company faces the following constraints

Minimum Production Demand (Units) 1,000 1,000
Machine Hours (MH) 5,000
Direct Labor Hours 4,000

What Sales mix will optimize profits?

What is the profit at this level of sales?

Use Solver for Constraints. Please show each step
Answer #1

The model is shown below

x fc B =121-122 C D Standard Premium 400 540 E F I Total J K L M N O P Q G H Standard Premium 0 Ε Mix 40 20 1.2 1.5 2 1 0 25

The formulas shown below

123 : x fc =121-122 A H Premium Standard Premium Standard Total 400 540 Mix 10 20 =C4*G$2 =C5*G$2 =C6*G$2 =D4*H$2 =D5*H$2 =D6

The solver parameters shown below

123 fr =121-122 Solver Parameters E F D Standard Premium 400 540 G H Standard Premium 0 0 Total Set Objective: SI523 Mix To:

The result is shown below. Optimum sales mix is 1665 standard and 1501 premium units. The maximum profit is 668550.5.

123 1 x A B fc =121-122 C D Standard Premium 400 540 E F L M N O P Q A G H Standard Premium 1665 1501 I Total Mix Solver Resu